Friday, October 20, 2006
- The testing of electrical panels on the upper floors at 1542 Tulane continued throughout the day.
- The small generator that had been powering temporary working lights in the basement was removed today. The lights are now being powered via the switch gear.
- Entergy requested that the test of the original VFD and filter be delayed until Monday because neither they nor the sub-contractor had sufficient staff available to man the effort today.
- The Student Parking Lot Lighting contractor finished running the conduit in the trenches. They are correcting one problem area where the conduit was clogged with dirt.
- At the Dental School, the Phase 2 mechanical contractor was on site making a check-list of all the items that require repair/replacement in the air handling unit rooms in both the Clinic and Administration Buildings .